# TAMO:Fine-Grained Root Cause Analysis via Tool-Assisted LLM Agent with Multi-Modality Observation Data

### v2

@@ IV-E RQ3: Ablation Study
  • $\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$: To investigate the effectiveness of the diffusion method used by tool $\mathscr{T}_{1}$ in fusing multimodal feature, in this variant we removed the tool $\mathscr{T}_{1}$ and directly used the raw time series data as input to evaluate the variant performance.
  • $\operatorname{TAMO}_{w/o\;branch}$: To investigate the effectiveness of the dual-branch diffusion model in tool $\mathscr{T}_{1}$, we removed the branch $\epsilon_{time}$, which is responsible for maintaining temporal features, in this variant. Only the log branch $\epsilon_{log}$ was used in the diffusion reconstruction process. The reconstructed data were then used for fault classification to evaluate variant performance.
  • $\operatorname{TAMO}_{w/o\;FFT}$: To investigate the effectiveness of the frequency domain-based root cause localization method proposed for tool $\mathscr{T}_{2}$, we removed the Fourier process in this variant. Instead, the original time-domain data were directly used for subsequent causal graph generation and graph convolution processes. The root cause localization results were then output to evaluate variant performance.
@@
Table II presents the results of the ablation experiments, from which it is evident that the removal of any component in TAMO leads to performance degradation. This demonstrates that each of the component within TAMO has an important effect. In particular, the variant of TAMO without tool $\mathscr{T}_{1}$ shows a significant decrease in performance. This is because the fact that the variant $\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$ lacks the ability of the diffusion model to fuse multimodal information, thereby being reduced to learning with only a single modal feature. This absence of key feature information makes accurate root cause analysis challenging. Similarly, removing the time branch $\epsilon_{time}$ led to noticeable decreases in performance for both tasks, indicating that time series features are crucial for accurate fault localization and classification. Eliminating the FFT process resulted in a substantial degradation in RE localization metrics, demonstrating that frequency domain analysis is essential for precise root cause localization.
IV-F RQ4: Case Study
@@
Subsequently, we input the model inference results along with system context information into the RCA expert LLM and compare its output with that of an LLM fed directly with raw data, as shown in Fig.5. This framework leverages tool information and domain expertise to accurately analyze current system faults and provide reasonable remediation suggestions. In contrast, the normal LLM using only raw data struggles to handle the vast amount of observational data due to severe context limitations, resulting in significant omissions and misjudgments in root cause identification and fault classification, as well as corresponding errors in the proposed remediation suggestions.
@@ V Conclusion
This paper introduces an innovative tool-assisted LLM agent framework named TAMO to address the complex challenges of root cause analysis in cloud-native microservice environments. By integrating domain-specific tools with large language models, TAMO efficiently handles high-dimensional, multi-modal cloud-native observability data and accurately locates root causes of failures in dynamically changing service dependencies. The approach not only overcomes the information loss issues when converting raw observability data into a format suitable for LLM processing but also addresses the challenge that static pre-trained knowledge struggles to adapt to real-time changes in service dependencies. The experimental results show that, compared to existing advanced methods and benchmarks, TAMO performs notably better on two standard datasets, highlighting its potential to improve fault diagnosis efficiency and accuracy. However, as system scale and complexity continue to grow, how to more efficiently process larger volumes of data streams and further enhance real-time interpretability and cross-modal semantic alignment remain critical areas for future research.

### v4

@@ IV-E RQ3: Ablation Study
  • $\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$: To investigate the effectiveness of the diffusion method used by tool $\mathscr{T}_{1}$ in fusing multimodal feature, in this variant we remove the tool $\mathscr{T}_{1}$ and directly use the raw time series data as input to evaluate the variant performance.
  • $\operatorname{TAMO}_{w/o\;branch}$: To investigate the effectiveness of the dual-branch diffusion model in tool $\mathscr{T}_{1}$, we remove the branch $\epsilon_{time}$, which is responsible for maintaining temporal features, in this variant. Only the log branch $\epsilon_{log}$ is used in the diffusion reconstruction process. The reconstructed data are then used for fault classification to evaluate variant performance.
  • $\operatorname{TAMO}_{w/o\;FFT}$: To investigate the effectiveness of the frequency domain-based root cause localization method proposed for tool $\mathscr{T}_{2}$, we remove the Fourier process in this variant. Instead, the original time-domain data are directly used for subsequent causal graph generation and graph convolution processes. The root cause localization results are then used to evaluate variant performance.
Table III presents the results of the ablation experiments, from which it is evident that the removal of any component in TAMO leads to performance degradation. This demonstrates that each of the component within TAMO has an important effect. In particular, the variant of TAMO without tool $\mathscr{T}_{1}$ shows a significant decrease in performance. This is because the fact that the variant $\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$ lacks the ability of the diffusion model to fuse multi-modality information, thereby being reduced to learn with only a single modal feature. The absence of key feature information makes accurate root cause analysis challenging. Similarly, removing the time branch $\epsilon_{time}$ lead to noticeable decreases in performance for both tasks, indicating that time series features are crucial for accurate fault localization and classification. Eliminating the FFT process result in a substantial degradation in RE localization metrics, demonstrating that frequency domain analysis is essential for precise root cause localization.
IV-F RQ4: Performance Study
To evaluate computational efficiency, we measure the training and inference time of TAMO on both datasets, as shown in Figure 5, with Eadro as a representative baseline. For training, TAMO requires 0.81 seconds per epoch on Dataset B and 5.7 seconds on Dataset A, indicating favorable scalability with minimal computational overhead relative to Eadro. For inference, TAMO processes each sample in 0.17 seconds, which is slightly slower than Eadro but still meets the real-time requirement for root cause analysis.
In addition, we provide the parameter counts for each component of the TAMO framework in Table IV. The multi-modality alignment tool $\mathscr{T}_{1}$ accounts for the majority of parameters (9.32M), as aligning heterogeneous multimodal data into a unified temporal representation requires a large number of parameters to learn effective alignment. In contrast, the root cause localization tool $\mathscr{T}_{2}$ and fault types classification tool $\mathscr{T}_{3}$ are significantly more lightweight (1.38M and 258.32K parameters, respectively).
TABLE IV: Parameter counts of the tool models in TAMO.
Model Parameters
Multi-modality Alignment Tool ($\mathscr{T}_{1}$) 9.32M
Root Cause Localization Tool ($\mathscr{T}_{2}$) 1.38M
Fault types Classification Tool ($\mathscr{T}_{3}$) 258.32K
Total parameters 10.96M
Finally, we analyze sensitivity to key hyperparameters (batch size, $\beta$, $\mu$, hidden dimension). As shown in Fig. 4, performance is highly sensitive to $\beta$, with minor changes causing sharp accuracy drops. In contrast, smaller batch sizes consistently yield better results, while $\mu$ causes significant performance degradation at its boundary values ($\mu$ = 0 or 1), indicating that it is designed to achieve critical branch balancing.
@@ IV-F RQ4: Performance Study
IV-G RQ5: Case Study
@@
Subsequently, we input the model inference results along with system context information into the RCA expert LLM and compare its output with that of an LLM fed directly with raw data, as shown in Fig.7. This framework leverages tool information and domain expertise to accurately analyze current system faults and provide reasonable remediation suggestions. In contrast, the normal LLM using only raw data struggles to handle the vast amount of observation data due to severe context limitations, resulting in significant omissions and misjudgments in root cause identification and fault classification, as well as corresponding errors in the proposed remediation suggestions.
@@ V Conclusion
In this paper, we introduced a tool-assisted LLM agent framework named TAMO to address the complex challenges of root cause analysis in cloud-native systems. By integrating domain-specific tools with large language models, TAMO efficiently handled high-dimensional, multi-modal cloud-native observation data and accurately located root causes of failures in dynamically changing service dependencies. The approach not only overcame the information loss issues incurred when converting raw observability data into a format suitable for LLM processing, but also addressed the challenge that static pre-trained knowledge struggled to adapt to real-time changes in service dependencies. The experimental results show that, compared to existing advanced methods and benchmarks, TAMO performs notably better on two standard datasets, highlighting its potential to improve fault diagnosis efficiency and accuracy.